3p(q-r^2) p=-1
q=-3
r=5
What is the solution? I got -66, but in class the teacher says its 84.
3(-1)(-3-5^2)
-3(-3+25)
-3(22)
-66


How do you get 84?

Answer by jim_thompson5910(35256)

Your work

3(-1)(-3-5^2)
-3(-3+25) <<<<------ It should be -25 since -5^2 = -25
-3(22)
-66


=======================================================


Corrected work

3p(q-r^2)

3(-1)(-3-5^2)

3(-1)(-3-25)

3(-1)(-28)

-3(-28)

84


So 3p(q-r^2) = 84 when p = -1, q = -3 and r = 5